### v3

@@ 1 Introduction
Figure 1: Different queries should trigger different reasoning modes. Simply cropping regions of interest and curating visual chain-of-thought reasoning is not a panacea.
@@ 2.3 Global Controller Module
Figure 3: Resolving the conflicts via V-RAG (chu2025reducing).
@@ 3.3 Ablation Study
Figure 6: figure
Effect of model scale on Orchestrator dispatch success and Synthesizer conflict-resolution rates.
Figure 7: figure
Ablation study for V-RAG, showing that larger $k$ increases the helpful retrieval rate but also raises the harmful rate.
@@ 4 Discussion
RadAgents leverages radiologist-inspired workflows, tool-augmented reasoning, and visual retrieval to improve CXR interpretation across three complex benchmarks. However, the system also depends on several external tools, making it sensitive to their errors or distribution shifts. Finally, our evaluation is restricted to offline datasets and automatic metrics; broader human studies and prospective assessments will be needed to understand clinical impact.
@@ Appendix C Vanilla-MedRAX-comparable experiments using GPT-4o
C.1 Experimental Setup
@@ C.2 Results
Table 2 summarizes performance across the three evaluation settings. RadAgents with V-RAG achieves the best results on all tasks. Relative to the ablated agent without V-RAG, it improves GREEN on VQA (E&A) from 0.5841 to 0.6032 (+0.0191, $\sim$3.3%), raises VQA (C&P) accuracy from 48.0% to 50.9% (+2.9 points, $\sim$6.0%), and boosts report-generation GREEN from 0.3821 to 0.4527 (+0.0706, $\sim$18.5%). Augmenting GPT-4o with ReAct and our workflow also yields consistent gains over the base LLM—for example, GREEN on VQA (E&A) increases from 0.2127 (GPT-4o) to 0.4619 (GPT-4o+ReAct) and 0.5351 (GPT-4o+Workflow), while VQA (C&P) accuracy rises from 18.2% to 41.9% and 45.5%, respectively. Nonetheless, even the strongest GPT-4o+Workflow baseline remains substantially behind RadAgents, trailing by 0.0681 GREEN on VQA (E&A), 5.4 accuracy points on VQA (C&P), and 0.0686 GREEN on report generation. Overall, the ranking of methods is largely consistent across tasks, suggesting that the benefits of structured tool use and V-RAG transfer from explanation-style VQA to longitudinal comparison questions and free-text report generation.

### v4

@@ 1 Introduction
Figure 1: Prior “encode-once, text-only” (left) and grounding/cropping-only variants (right) can fail on queries requiring iterative re-inspection and quantitative assessment (e.g., cardiothoracic ratio). RadAgents instead supports multimodal interleaved reasoning, where hypotheses trigger targeted visual operations and the final answer is grounded in explicit visual evidence.
@@ 2.3 Global Controller Module
Figure 3: V-RAG Mechanism. When tool outputs conflict, the system retrieves top-$k$ clinically similar CXR studies as reference standards to verify the findings and resolve the disagreement (e.g., Edema and Cardiomegaly).
@@ 3.3 Ablation Study
Figure 6: Effect of model scale on Orchestrator dispatch success and Synthesizer conflict-resolution rates.
Figure 7: Pilot study for $k$ selection in V-RAG, showing that larger $k$ increases the helpful retrieval rate but also raises the harmful rate.
@@ 4 Discussion
RadAgents leverages radiologist-inspired workflows, tool-augmented reasoning, and visual retrieval to achieve robust CXR interpretation across three complex benchmarks. Despite these advancements, the current framework has limitations. First, performance is intrinsically upper-bounded by the capabilities of the underlying tools; specifically, the current reliance on tools optimized for frontal-view X-rays limits robustness on lateral views. Second, the interaction between the orchestrator and tools is unidirectional: while the synthesizer can resolve conflicts via evidence weighting, it cannot iteratively guide or correct upstream tool outputs (e.g., refining an imperfect segmentation mask). Additionally, the multi-agent architecture incurs higher computational costs compared to end-to-end baselines. Future work will address these challenges by optimizing efficiency through dynamic agent selection and extending the framework to 3D modalities (e.g., CT and MRI) via modality-specific sub-agents, alongside prospective studies to validate clinical impact.
@@ Appendix C Vanilla-MedRAX-comparable experiments using GPT-4o
C.1 Implementation Details of RadAgents
C.2 Experimental Setup
@@ C.3 Results
Table 2 summarizes performance across the three evaluation settings. RadAgents with V-RAG achieves the best results on all tasks. Relative to the ablated agent without V-RAG, it improves GREEN on VQA (E&A) from 0.5841 to 0.6032 (+0.0191, $\sim$3.3%), raises VQA (C&P) accuracy from 48.0% to 50.9% (+2.9 points, $\sim$6.0%), and boosts report-generation GREEN from 0.3821 to 0.4527 (+0.0706, $\sim$18.5%). Augmenting GPT-4o with ReAct and our workflow also yields consistent gains over the base LLM; for example, GREEN on VQA (E&A) increases from 0.2127 (GPT-4o) to 0.4619 (GPT-4o+ReAct) and 0.5351 (GPT-4o+Workflow), while VQA (C&P) accuracy rises from 18.2% to 41.9% and 45.5%, respectively. Nonetheless, even the strongest GPT-4o+Workflow baseline remains substantially behind RadAgents, trailing by 0.0681 GREEN on VQA (E&A), 5.4 accuracy points on VQA (C&P), and 0.0686 GREEN on report generation. Overall, the ranking of methods is largely consistent across tasks, suggesting that the benefits of structured tool use and V-RAG transfer from explanation-style VQA to longitudinal comparison questions and free-text report generation.
@@ Appendix D Details of V-RAG
Corpus.
The retrieval index consists exclusively of the MIMIC-CXR official training split. We strictly enforce patient-level exclusion (via Subject ID), ensuring test patients are completely disjoint from the retrieval corpus. For external benchmarks (e.g., ChestAgentBench), the data originates from distinct sources, ensuring zero overlap with the MIMIC-CXR index.
@@
Appendix E Additional Results and Analysis
E.1 Entity-level Evaluation for Report Generation
Evaluating the quality of generated radiology reports is non-trivial. Early works adopted general-domain natural language processing metrics such as ROUGE (lin2004rouge) and BLEU (papineni2002bleu). While these metrics are widely used for text evaluation, they treat differences in wording the same as clinically significant errors, failing to reflect medical accuracy. To address this limitation, clinically informed evaluation metrics, such as CheXbert (smit2020combining), RadGraph (jain1radgraph), GREEN (ostmeier2024green), and RaTEScore (zhao2024ratescore), have been proposed to better assess clinical correctness and utility. CheXbert is based on multi-label classification results for 5 or 13 diseases (along with one extra “normal” label). RadGraph considers literal entity agreement considering the positive or negative context of each entity. GREEN judges recall and precision errors by LLM prompting. RaTEScore is inspired by RadGraph but less sensitive to phrasing by an F1-like computation which allows semantic matching between entities based on a cosine similarity. The metrics are computed using their official and standardized implementations: RadGraph-F1444https://pypi.org/project/radgraph/0.1.2/, CheXbert-F1555https://pypi.org/project/f1chexbert/, RaTE Score [^6] https://pypi.org/project/RaTEScore/0.5.0/, and GREEN [^7] https://pypi.org/project/green-score/0.0.8/.
Table 3: Performance Comparison on Report Generation Metrics. The proposed RadAgents achieves the highest scores across all standard metrics.
Settings / Metrics CheXbert-macro-F1 (14) RadGraph-F1 RaTE
CheXagent 29.2 12.7 44.4
GPT-4o 22.4 15.3 49.8
Qwen3-VL w/ ReAct 44.9 17.1 53.6
Qwen3-VL w/ Workflow 45.5 17.2 53.9
RadAgents w/o V-RAG 49.1 18.1 55.8
RadAgents (Ours) 53.2 19.4 58.5
E.2 Computational Cost and Latency Analysis
To assess the deployment feasibility of RadAgents, we conducted a detailed analysis of inference latency and computational cost.
Experimental Setup and Mechanism.
A critical efficiency feature of RadAgents is its sparse activation design. As detailed in Section 2.3, the Orchestrator activates strictly those sub-agents required for a specific query (e.g., a query regarding “bone fractures” triggers only the relevant specialist, leaving the “Lung Opacity” agent idle). Consequently, the system rarely incurs the computational overhead of running all 7 agents simultaneously; idle agents consume zero active compute resources.
Performance was measured on an on-premise node equipped with $8\times$ NVIDIA RTX A5000 (24GB) GPUs. To reflect a practical deployment environment, we utilized quantized models served via Ollama backends with LangGraph orchestration. We report two key metrics: (1) Latency: End-to-end wall-clock time per case (seconds), and (2) Compute Cost: Cumulative GPU usage per case (GPU-seconds).
Quantitative Results.
We compared the proposed RadAgents (in both Sequential and Parallel execution modes) against single-agent baselines across three benchmarks. The results are summarized in Table 4.
Table 4: Latency and Cost Analysis across benchmarks. Metrics are reported as Wall-Clock Time (sec) / Compute Cost (GPU-sec). The Parallel configuration significantly reduces latency compared to sequential execution.
Configuration CheXbench ChestAgentBench Report Gen
(sec / GPU-sec) (sec / GPU-sec) (sec / GPU-sec)
Single-Agent ReAct 16 / 30 25 / 50 50 / 110
Single-Agent Workflow 18 / 36 29 / 64 58 / 140
RadAgents (Sequential) 38 / 89 62 / 178 120 / 368
RadAgents (Parallel) 26 / 83 41 / 164 85 / 340
Analysis.
Parallelism Efficiency: Running sub-agents in parallel reduces wall-clock latency by approximately 30–40% compared to the sequential variant. For complex queries in ChestAgentBench, the system achieves an average latency of 41 seconds, rendering it responsive enough for asynchronous clinical workflows.
[…]
E.3 Systematic Error Analysis
To evaluate the robustness of the Orchestrator/Sub-agents and the stability of the agentic workflows, we analyzed the frequency of workflow deviations across three benchmarks. Specifically, we tracked two key metrics:
  1. 1.
    Workflow-free (ReAct) Fallback Rate: The percentage of queries where the Orchestrator could not match a pre-defined clinical workflow and defaulted to a generalist ReAct loop.
  2. 2.
    SkillMismatchError / Re-dispatch Rate: The frequency with which an assigned sub-agent rejected a task (due to scope mismatch) or failed, necessitating a re-dispatch by the Orchestrator.
Quantitative Results.
The average error frequencies are summarized in Table 5.
Table 5: Systematic Error Frequency Analysis. ChestAgentBench exhibits higher fallback rates due to its open-ended query nature, whereas MIMIC-CXR report generation follows a highly structured routine, resulting in minimal deviations.
Datasets Workflow-free (ReAct) SkillMismatchError /
Fallback Rate Re-dispatch Rate
CheXbench 10.4% 5.3%
ChestAgentBench 18.1% 9.3%
MIMIC-CXR Report Gen 2.1% 4.0%
Interpretation.
The variance in error rates reflects the distinct nature of the evaluation datasets:
  • ChestAgentBench contains a higher proportion of open-ended, mixed-intent queries. This complexity forces the Orchestrator to utilize the fallback ReAct mechanism more frequently (18.1%) and results in higher rates of internal re-dispatching (9.3%) to resolve ambiguities.
  • MIMIC-CXR Report Generation typically follows a fixed, routine clinical workflow (e.g., standard frontal/lateral review). Consequently, it exhibits the highest stability with a minimal workflow-free fallback rate (2.1%).
  • CheXbench falls between these two extremes, representing a balanced mix of structured classification tasks and moderately complex reasoning queries.
Appendix F Demonstration
An end-to-end execution trace for the query “Is the trachea midline?”. (a) Dispatch: The Orchestrator routes the query to the Airway Agent based on intent analysis. (b) Tool Execution: The sub-agent invokes the segmentation tool; note the incorrect ”midline” classification in the raw tool metadata despite successful segmentation. (c) Conflict Detection: The Synthesizer performs a cross-modal check using a VQA model, detecting a discrepancy between the tool’s heuristic and the visual assessment. (d) Resolution: The Synthesizer triggers V-RAG to retrieve Top-3 similar cases. By synthesizing the retrieved evidence with its own reasoning, the system corrects the tool error to output the final accurate diagnosis of “deviated to the right.”
Execution Trace: Query “Is the trachea midline?” STEP 1: ORCHESTRATION & DISPATCH   [ORCHESTRATOR] Analyzing Query Context...
Query: ‘‘Is the trachea midline?’’ Intent: Anatomic Localization (Airway)
Strategy: Activate [Airway_Agent] Dispatch: > Executing Airway_Agent...
STEP 2: SUB-AGENT REASONING & TOOL EXECUTION   [AIRWAY_AGENT] Action: Calling tool ‘check_tracheal_position‘ Params: { ‘‘organ’’: ‘‘Trachea’’ } [TOOL_OUTPUT] Status: SUCCESS Metrics: { ‘‘area_pixels’’: 12490, ‘‘centroid’’: [157.8, 470.0], ‘‘heuristic_classification’’: ‘‘midline’’ <-- (Incorrect) } [AIRWAY_AGENT] Tool execution complete. > Handoff to Synthesizer. [Uncaptioned image] Tool Segmentation Visualization STEP 3: CONFLICT DETECTION (SYNTHESIZER)   [SYNTHESIZER] Phase 1: Verification
> Tool Claim: ‘‘Trachea is midline’’ (based on centroid heuristic).
> Action: Cross-check with VQA (Global View). > VQA: ‘‘The trachea appears deviated to the right.’’
! ALERT: Tool output contradicts visual assessment. ! Status: UNCERTAIN. Triggering V-RAG Resolution... STEP 4: RESOLUTION (V-RAG + SYNTHESIS)   [SYNTHESIZER] Phase 2: Resolution via V-RAG
> Action: Retrieving Top-3 clinically similar cases...
[RETRIEVAL] Found 3 references (Sim >> 0.85):
1. Case_782: ‘‘Trachea deviated to right’’ (Visual match: High) 2. Case_104: ‘‘Rightward deviation due to effusion’’ (Visual match: High) 3. Case_993: ‘‘Midline trachea’’ (Visual match: Low) [FINAL REASONING]
‘‘Although segmentation suggests midline, VQA and 2/3 retrieved cases confirm rightward deviation. The tool’s centroid heuristic is likely misled by patient rotation.’’
[FINAL ANSWER] ‘‘No, the trachea is not midline; it is deviated to the right.’’ Confidence: High (Supported by V-RAG Evidence)